If log_x^2(9)-log_x(36)=1
What does x equal?

Answer by lwsshak3(11628)

If log_x^2(9)-log_x(36)=1
What does x equal?
**
log(9x^2)-log(36x)=1
log[(9x^2)/(36x)]=1
convert to exponential form
10^1=(9x^2)/(36x)=10
x/4=10
x=40